# High-Speed Event Vision-Based Tactile Roller Sensor for Large Surface Measurements

### v1

@@ V-C High-Speed Tactile Braille Reading
The large surface tactile scanning of the roller was also tested with regards to feature detection and recognition. Two braille plates were scanned with standard braille sizing: semi sphere of radius of 0.635 mm and spacing of 2.54 mm. The processing pipeline in Section III-F was utilized to process the events into a frame and subsequently detect the letter in the frame. Each frame is then stitched to the next frame to form the complete words (as seen in Figure 11 and Figure 1). Due to the high temporal resolution (microsecond resolution) of the event camera, the recognition algorithm performs well at low and high speeds (0.05 to 0.5 m/s). Utilizing the formula presented by Potdar et al. [23] for converting from m/s to wpm, our roller performs braille reading at 831 wpm without loss in accuracy, which is 2.6 times faster than the speed they demonstrate at 87.5% accuracy. Furthermore, as Potdar et al. highlight, this tactile feature recognition pipeline can be extended for dynamic detection of surface textures [23].
@@
Figure 11: High-speed braille reading. Tactile braille plate (top) and detected letters grid (bottom).
VI Conclusion
In this work, we introduced a novel high-speed neuromorphic vision-based roller tactile sensor designed for continuous large surface inspection, addressing the speed limitations inherent in conventional VBTS technologies. We presented the sensor’s mechanical design, integrating an event camera within a rolling elastomer mechanism, and developed a 3D reconstruction pipeline. This pipeline adapts the event-based multi-view stereo (EMVS) method and incorporates a Bayesian Model Averaging (BMA) fusion strategy, combining depth information from multiple reference viewpoints within a temporal window to enhance accuracy and robustness.
@@ VI Conclusion
Our experimental results demonstrate the significant capabilities of the proposed system. We achieved state-of-the-art continuous tactile scanning speeds of up to 0.5 m/s for 3D surface reconstruction, yielding Mean Absolute Errors (MAE) below 100 microns on large surface tasks. This speed represents a substantial leap, being approximately 11 times faster than reported belt-based sensors and 45 times faster than previous roller-based VBTSs. The BMA fusion technique proved crucial, reducing reconstruction MAE by an average of 25.2% compared to standard EMVS. Furthermore, leveraging the sensor’s high temporal resolution, we demonstrated Braille character recognition at 831 words per minute, 2.6 times faster than the current benchmark for tactile Braille reading. These findings underscore the potential of neuromorphic vision to revolutionize high-speed tactile inspection, enabling applications previously infeasible due to the speed constraints of traditional methods, particularly in large-scale industrial quality control for aerospace or automotive components.
@@
Future research will focus on addressing these limitations and expanding the sensor’s capabilities. Firstly, we plan to investigate the integration of force and contact angle sensing and control. This could involve adding markers to the non-sensing sides of the roller, using their displacement within a learning-based framework to estimate applied force and roller orientation relative to the surface. Implementing closed-loop control based on this feedback would ensure consistent contact, which is crucial for maintaining high-quality tactile data acquisition, especially on non-planar surfaces. Secondly, to overcome the challenges associated with sparse depth maps from EMVS, we aim to explore alternative event-based reconstruction methods capable of generating denser surface information. Techniques such as event-based structured light measurement or event-based photometric stereo could potentially be adapted to the tactile roller geometry to directly estimate dense surface normals or depth maps from the event data. Achieving denser reconstructions would significantly enhance the sensor’s ability to accurately measure complex, continuously varying surface geometries, broadening its applicability.

### v2

@@ VI-C High-Speed Tactile Braille Reading
The processing pipeline above was utilized to process the events into a frame and subsequently detect the letter in the frame. Each frame is then stitched to the next frame to form the complete words (as seen in Figure 1 and Figure 13). Due to the high temporal resolution (microsecond resolution) of the event camera, the recognition algorithm performs well at low and high speeds (0.05 to 0.5 m/s). For comparison, Potdar et al. [potdar_high-speed_2024] demonstrated Braille reading at approximately 0.18 m/s with 87.5% accuracy. Our system achieves the same task at 0.5 m/s, which is 2.6 times faster, while maintaining full accuracy. Furthermore, as Potdar et al. highlight, this tactile feature recognition pipeline can be extended for dynamic detection of surface textures [potdar_high-speed_2024].
@@
Figure 13: High-speed braille reading. (a) Tactile braille plate (b) Close-up of a few detected braille letters (c) Full detected letters grid
VI-D Processing Time and Computational Efficiency
All experiments were run on a laptop CPU (13th Gen Intel Core i7-1365U) with offline processing of the recorded event data. The Braille reading pipeline, consisting of frame generation and Braille detection, achieved an average processing time of 19.26 ± 0.77 ms per frame (averaged over five runs). In comparison, the EMVS 3D reconstruction pipeline (from events to depth map) required 377.41 ± 7.96 ms per frame under the same conditions. This over 19× difference in processing time highlights the advantage of direct 2D feature recognition for tasks such as Braille reading, where the features of interest are small, well-defined, and do not require full 3D reconstruction for accurate detection. Although it’s not explored in this work, the low processing time for the Braille pipeline indicates its strong potential for real-time operation. Conversely, the current CPU-based EMVS implementation is not real-time, as its processing time exceeds the 20 ms data acquisition window that was set for these experiments. However, leveraging a more powerful CPU and GPU acceleration for the highly parallelizable ray-casting and voting process could significantly improve performance and enable real-time 3D reconstruction.
VII Discussion And Limitations
@@
Our experimental results demonstrate the significant capabilities of the proposed system. We achieved state-of-the-art continuous tactile scanning speeds of up to 0.5 m/s for 3D surface reconstruction, yielding Mean Absolute Errors (MAE) below 100 microns on large surface tasks. This speed represents a substantial leap, being approximately 11 times faster than reported belt-based sensors and 45 times faster than previous roller-based VBTSs. The BMA fusion technique proved crucial, reducing reconstruction MAE by an average of 25.2% compared to standard EMVS. Furthermore, leveraging the sensor’s high temporal resolution, we demonstrated Braille character recognition at 0.5 m/s, 2.6 times faster than the current benchmark for tactile Braille reading. These findings underscore the potential of neuromorphic vision to revolutionize high-speed tactile inspection, enabling applications previously infeasible due to the speed constraints of traditional methods, particularly in large-scale industrial quality control for aerospace or automotive components. This high-fidelity 3D reconstruction capability is crucial, as it allows for the direct comparison of a scanned surface against its reference CAD model to automatically detect and quantify volumetric defects.
While the reliance on distinct edges and sharp features is a limitation of the underlying EMVS algorithm, it is important to contextualize this within the scope of industrial surface inspection. The primary goal in such applications is often the detection and quantification of high-frequency geometric anomalies, such as cracks, scratches, dents, weld seams, or the precise profiles of fasteners and drilled holes. These critical defects are inherently defined by the sharp edges and distinct features that our event-based system is specialized to capture with high fidelity. Therefore, while our method may not be ideal for reconstructing smooth, featureless surfaces, its strength aligns directly with the requirements for identifying the most common and critical types of defects in industrial quality control.
An interesting observation from the calibration evaluation (Section V-B) is that the start reference view ($t_{s}$) consistently receives the highest weight, while the end view ($t_{e}$) receives the lowest. Intuitively, one might expect the midpoint view ($t_{m}$) to be most reliable, but several factors may contribute to this trend. First, the pose at $t_{s}$ is derived directly from the robot state at the start of the window, whereas $t_{m}$ and $t_{e}$ rely on interpolation and may accumulate small kinematic or timing errors. Second, as the roller moves, tactile features first appear and generate high-contrast events near $t_{s}$, potentially providing more stable features for EMVS before disocclusions occur. Finally, the DSI integrates events over the entire window, so errors later in the window may degrade the quality of the $t_{e}$ viewpoint more than the $t_{s}$ viewpoint. A more precise analysis of how the timing of the three reference windows affects reconstruction performance is also an interesting direction for future work.
@@
Future research will focus on addressing these limitations and expanding the sensor’s capabilities. Firstly, we plan to investigate the integration of force and contact angle sensing and control. This could involve adding markers to the non-sensing sides of the roller, using their displacement within a learning-based framework to estimate applied force and roller orientation relative to the surface. Implementing closed-loop control based on this feedback would ensure consistent contact, which is crucial for maintaining high-quality tactile data acquisition, especially on non-planar surfaces. Secondly, to overcome the challenges associated with sparse depth maps from EMVS, we aim to explore alternative event-based reconstruction methods capable of generating denser surface information. Techniques such as event-based structured light measurement or event-based photometric stereo could potentially be adapted to the tactile roller geometry to directly estimate dense surface normals or depth maps from the event data. Achieving denser reconstructions would significantly enhance the sensor’s ability to accurately measure complex, continuously varying surface geometries, broadening its applicability. Finally, a systematic investigation into the mechanical robustness and long-term durability of the elastomer is a crucial next step. Future work will involve quantifying the wear characteristics of the silicone skin over extended operational cycles and performing a mechanical analysis to determine the optimal elastomer parameters for roller-based applications
VIII Conclusion
In this work, we introduced a novel high-speed neuromorphic vision-based roller tactile sensor designed for continuous large surface inspection, addressing the speed limitations inherent in conventional VBTS technologies. We presented the sensor’s mechanical design, integrating an event camera within a rolling elastomer mechanism, and developed a 3D reconstruction pipeline. This pipeline adapts the event-based multi-view stereo (EMVS) method and incorporates a Bayesian Model Averaging (BMA) fusion strategy, combining depth information from multiple reference viewpoints within a temporal window to enhance accuracy and robustness. The proposed system demonstrates significant advances in tactile sensing speed and accuracy, achieving continuous scanning at 0.5 m/s with sub-100 micron precision, representing an order-of-magnitude improvement over existing continuous tactile sensing approaches. This breakthrough enables practical deployment for large-scale industrial surface inspection applications previously constrained by the speed limitations of conventional tactile sensors. By enabling rapid, high-fidelity 3D reconstruction of complex surfaces, this technology has the potential to significantly improve quality assurance workflows in industries such as aerospace, automotive, and precision manufacturing, reducing inspection times and increasing defect detection reliability.